OpenStudy (anonymous):

What is 40 km^3 -> m^3?

OpenStudy (anonymous):

someone answer this please? (With steps)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

40 km^3 = 40 * (1000m)^3 = 40 * (1000)^3m^3 = 40 * 10^9 m^3

OpenStudy (anonymous):

this is the right one

OpenStudy (anonymous):

so the answer is (40)(10^9m^3)?

OpenStudy (anonymous):

(40*10^9) m^3

OpenStudy (anonymous):

ahh, okay :) thank you!!